Angel scored a total of 80 points in a basketball game against Jason. Angel made a total of 37 two-point
and three-point baskets. How many two-point shots did Angel make? How many three-point shots did he make?

Answer by rothauserc(4718)

let x be the number of two-point baskets and y be the number of 3-point baskets
:
(1) x + y = 37
:
(2) 2x +3y = 80
:
solve equation 1 for x and substitute in equation 2
:
x = 37 -y
:
2(37-y) +3y = 80
:
74 -2y +3y = 80
:
y = 80 -74 = 6
:
x = 37 -6 = 31
:
*********************************************************
Angel made 31 two-point baskets and 6 three-point baskets
:
check answer by substituting x and y values in equation 2
:
2(31) +3(6) = 80
:
62 +18 = 80
:
80 = 80
:
answer checks
*********************************************************
: